1.the following costs were incurred by odd a mfg cost

DM =105200

DL = 221400

MOH = 98600

based on the above data

a.determine prime cost

b.conversion cost

c.mfg cost


1
Expert's answer
2021-08-18T11:05:28-0400

Solution:

a.). Prime cost = Direct Materials + Direct Labor

Direct Material used = 105,200

Direct Labor used = 221,400

Prime cost = 105,200 + 221,400 = 326,600

 

b.). Conversion cost = Direct labor + Manufacturing Overheads

Conversion cost = 221,400 + 98,600 = 320,000

 

c.). Manufacturing cost = Direct Materials + Direct Labor + Manufacturing Overheads

Manufacturing cost = 105,200 + 221,400 + 98,600 = 425,200